# Earth Observation Foundation Model PhilEO: Pretraining on the MajorTOM and FastTOM Datasets

### v2

@@ V-F Discussion of main findings, and per-class results
Per-class results for land cover mapping. We also evaluate and examine the precision metrics for each class across different $n$-shots settings in Fig. 21, focusing on the PhilEO 44M MajorTOM 23TB model’s performance on the PhilEO Bench land cover semantic segmentation task, which includes $11$ classes based on the ESA WorldCover classification scheme. Similarly, Fig. 22 presents the class-wise precision metrics for the PhilEO 44M FastTOM 2TB model on the same land cover semantic segmentation task. The results from both figures demonstrate that precision generally improves across most of the classes as the number of $n$-shot examples increases. At this point, we also note that the $11$ classes in the ESA WorldCover semantic segmentation dataset in Figs. 21 and 22 are: Tree cover, Shrubland, Grassland, Cropland, Built-up, Bare/sparse vegetation, Snow and ice, Permanent water bodies, Herbaceous wetland, Mangrove, and Moss and lichen.
@@ VI Conclusion
In this paper, we also compared with Mamba models (see Sec. III-C, as well as Sec. V-D), and this also one of our main contributions. Furthermore, we think our FLOPs profiling analysis, comparison, and benchmarking (i.e. in Sec. V, as well as in Figs. 8 and 9) opens the road to fairly evaluating different models (including GFMs) and will inspire others to adopt and use this real-world setting. Finally, it was a major achievement to process the 23TB of MajorTom on 32 cores of the Leonardo supercomputer davinci-1, in many days. This would have taken many more days, and even more months, on a single GPU. As future work, we plan to use JEPA, as well as add multi-temporal capability (i.e. multiple time steps) [58], and multi-modality. Regarding the latter, we will add S-1 in addition to S-2 in the pretraining and not only in the S12-PhilEOBench (i.e. see Sec. III-D). As future work, we will also use Mamba 2D models with larger images, for example 512512x512512 (or even 10241024x10241024), as well as perform knowledge distillation of models to the same number of FLOPs (e.g., for On-Board AI) and compare and benchmark. Finally, as future work, we plan to also examine, compare with, and try to include in the model pipeline the recent AlphaEarth Foundations Satellite Embeddings. These are geo-located embeddings at 1010m resolution, and they also need the longitude and latitude information to be used in the downstream tasks.

### v3

@@ V-F Discussion of main findings, and per-class results
Per-class results for land cover mapping. We also evaluate and examine the precision for each class across different $n$-shot settings in Fig. 21, focusing on the PhilEO 44M MajorTOM 23TB model’s performance on the PhilEO Bench land cover semantic segmentation task, which includes $11$ classes based on the ESA WorldCover classification. Similarly, Fig. 22 presents the class-wise precision metrics for the PhilEO 44M FastTOM 2TB model on the same land cover semantic segmentation task. The results from the figures demonstrate that precision improves across most of the classes as the number of $n$-shots increases. Here, we also note that the $11$ classes in the WorldCover semantic segmentation dataset in Figs. 21 and 22 are: Tree cover, Shrubland, Grassland, Cropland, Built-up, Bare/sparse vegetation, Snow and ice, Permanent water bodies, Herbaceous wetland, Mangrove, and Moss and lichen.
Summary and final discussion. We have examined in this paper the effect of scaling models and data: i) Model size and architecture, and ii) Data size. The main reason for wide scaling is to include oceans and ice in addition to land, and we note that we defer for future work ocean downstream tasks, i.e. MADOS dataset [66] for marine litter detection. In this work, the main research questions (which we answer affirmatively to both) are: “Is it feasible to have a model tackling a wider variety of downstream tasks (i.e. oceans and ice) compared to smaller models?”, and “Does the extra information in MajorTOM 23TB (oceans, ice) make it possible to use the model?”. For the former research question, we note that the larger model is able to tackle a wider variety of downstream tasks. Our focus is on examining if the downstream tasks can be performed at the same level (or better) with a larger model.
@@ VI Conclusion
In this paper, we also compared with Mamba models (see Sec. III-C, as well as Sec. V-D), and this also one of our main contributions. Furthermore, we think our FLOPs profiling analysis, comparison, and benchmarking (i.e. in Sec. V, as well as in Figs. 8 and 9) opens the road to fairly evaluating different models (including GFMs) and will inspire others to adopt and use this real-world setting. Finally, it was a major achievement to process the 23TB of MajorTom on 32 cores of the Leonardo supercomputer davinci-1, in many days. This would have taken many more days, and even more months, on a single GPU. As future work, we plan to use JEPA, as well as add multi-temporal capability (i.e. multiple time steps) [58], and multi-modality. Regarding the latter, we will add S-1 in addition to S-2 in the pretraining and not only in the S12-PhilEOBench (i.e. see Sec. III-D). As future work, we will also use Mamba 2D models with larger images, for example 512512x512512 (or even 10241024x10241024), as well as perform knowledge distillation of models to the same number of FLOPs (e.g., for On-Board AI) and compare and benchmark. Finally, as future work, we plan to also examine, compare with, and try to include in the model pipeline the recent AlphaEarth Foundations Satellite Embeddings [^8] http://deepmind.google/discover/blog/alphaearth-foundations-helps-map-our-planet-in-unprecedented-detail/. These are geo-located embeddings at 1010m resolution, and they also need the longitude and latitude information to be used in the downstream tasks.